OpenStudy (anonymous):

Please Help!!! The dependence of the equilibrium constant of a reaction on temperature is given by the van’t Hoff equation: Ink=-deltaH^0/RT+C Therefore, the ratio of the equilibrium constants at two different temperatures is: In K^1/K^2=deltaH^0/R(1/T2-1/T1) Calculate the pH of water at its normal boiling point. Is the water acidic, basic, or neutral?

OpenStudy (paxpolaris):

\[\LARGE \ln \left( K_1\over K_2 \right)={\Delta H^\Theta \over R}\left(\frac 1 T_2-\frac 1 T_1\right)\]

OpenStudy (anonymous):

thats the question... do you know the answer?? and hw did you did you do thatt??

OpenStudy (anonymous):

thats the question... do you know the answer?? and hw did you did you do thatt??

OpenStudy (paxpolaris):

Do you know ΔH^Θ for water ?

OpenStudy (anonymous):

No I cant find itt in the book that is why my friennd told me to come on this website and someone will help me me... so can you?

OpenStudy (paxpolaris):

I'm just guessing here: for the reaction: \(\large H_2O \rightarrow H^+ +\ OH^-\) \[\Large \Delta H^\Theta= \left( \Large \Delta H^\Theta_{f (H^+)} +H^\Theta_{f (OH^-)}\right)-H^\Theta_{f (H_2O)}\] \[\Large = (0+(-229.9))-(-289.8)=55.9{kJ \over mol}\]

OpenStudy (anonymous):

Thank you